# Learning Object Placement Programs for Indoor Scene Synthesis with Iterative Self Training

### v1

@@ 5.2. Object Encoding
The object encoder encodes object bounding boxes into an embedding vector $\in\mathbb{R}^{d}$. A learned embedding of the object category is concatenated to the raw values of the other attributes and passed through an MLP. Instead of encoding the floor plan, and implicitly the walls, as a single object, we encode each wall segment as its own object. This is due to the difficulty of encoding all the pertinent information about a floor plan into a single feature vector, and giving the model access to that information in a salient manner. The model struggles to infer algebraic quantities such as object to wall distances, especially for floor plans with non-convex geometry.
5.3. Program Encoding and Decoding
We represent programs as two separate sequences. One sequence represents the topology or structure of the program tree, and the other represents the attributes of each constraint in the program. We flatten out the program’s tree structure with prefix notation. We embed this “structure” sequence with per token learnable embeddings.
@@ 5.3. Program Encoding and Decoding
Each constraint takes the form (constraint type, query object index, reference object index, direction) or $(c_{j},q_{j},r_{j},d_{j})$. Constraint attributes are concatenated into a sequence following inorder traversal (left subtree to root to right subtree). The constraint type and direction receive per token learnable embeddings. Tokens which represent the query or reference object index use their respective object embeddings generated by the object encoder.
@@
For tokens of fixed vocabulary length such as program structure, constraint type, and direction, an MLP head is enough to decode their tokens, but when choosing the reference object index the vocabulary is variable length as the number of objects in a scene is variable. To address this, we pass each reference object head through an MLP to form a pointer embedding $v_{j}\in\mathbb{R}^{d}$(Vinyals et al., 2015). For a matrix of object embeddings $X\in\mathbb{R}^{txd}$ where $t$ is the number of objects in the scene, we compute the reference object index $r_{j}$ as
(1) $r_{j}=\text{arg max}(\text{Softmax}(Xv_{j}))$
@@
The dot product of the pointer embedding with the object embeddings forms a probability distribution over the objects. The reference object is the object with the highest probability mass. S
6. Program Self Training
@@
In this section we describe our program self training algorithm which discovers DSL programs from 3D scene data. It also helps improve next-object distribution locations predicted by our system. An overview of our algorithm is shown in Figure 3.
@@
Our algorithm falls under the PLAD (Jones et al., 2022) family, a conceptual framework for unsupervised program bootstrapping. These methods iteratively improve a dataset of programs by searching for new and better programs, retraining on them, and then repeating the process. Our algorithm also takes inspiration from Talton et al. (Talton et al., 2012), a work which uses probabilistic context free grammars (PCFGs) to learn a procedural model from a set of examples. The optimization begins with the “most-specific” grammars and converges to a grammar which is not too specific and not too general.
@@
Our optimization begins with programs that represent a subset of all possible valid placements. For each object in a scene, we use geometric heuristics to apply every possible constraint to the object so that the extracted program will only place the object where it was originally found. Further details are described in the supplemental material. We add additional placement modes to these “most restrictive” programs through a search and filtering process. Similar to Ganeshan et al. (Ganeshan et al., 2023) we use domain specific operations to edit a subset of programs per iteration of self training. This is both for computational feasibility and distributional stability between iterations.
6.1. Candidate Program Generation
@@
For a given partial scene and object to add we search for new programs by sampling the generative model described in Section 5 and then relaxing both the inferred and original program. Program relaxation for our representation entails randomly removing constraints from the program tree to produce a new program. This can help generalize the overly restrictive programs produced by our initial naive approach. For example, an object found in the corner of a room might initially be constrained to rest against both walls. Removing one of these constraints would allow the program to predict more general placements along one of the walls.
@@
We employ simple filtering techniques on candidate programs. Severely unconstrained programs that elect to place the object anywhere in the scene and programs which do not predict any placements are removed from consideration. We also ensure that each candidate program predicts only one valid orientation. Programs which predict placements for multiple orientations are split into subtrees that each predict only one. This constraint reduces the number of repeated subtrees present in the final programs and improves quantitative performance of this algorithm. It is worth noting that this constraint does not restrict the possible orientations the final program can represent. Our later program combination step produces final programs which predict valid placements for multiple different object orientations.
6.2. Scene Classifier
@@
While it might seem unintuitive to use a scene classifier to classify programs, we choose this route because both hard positives and negatives for programs are hard to generate. We train a real fake scene classifier to predict how in or out of distribution a particular object placement is. Positive examples are generated by randomly subsampling scenes. Negative examples come from randomly perturbing the rotation and location of a single object. Even though this process can generate negative examples that should be labeled positive, we find it sufficient for learning a useful decision boundary. We sample a program’s predicted mask multiple times, insert the object in question at those sampled locations, and then compute the scene classifier’s real probability for each insertion. The program’s final score is the average of these probabilities.
We perform experiments on two architectures of scene classifiers. We test a CNN based classifier that uses the input representation of Ritchie et al. (Ritchie et al., 2018) with an additional input channel denoting the query object in the scene. We also test a transformer based classifier that uses the object encoder described in section 5 with an additional learned vector embedding added to the query object vector encoding. While our CNN based classifier reports higher precision (accepts less invalid programs), and results in better downstream quantitative metrics we report results using the transformer classifier due to constraints on computation. The transformer model is computationally cheaper than the CNN because it does not require rasterization of the top down view of the scene. More details on the scene classifiers and their evaluation are included in the supplemental material.
6.3. Combining Programs
@@
Recall that our programs takes on the structure of a CSG tree. If two candidate programs predict two different placement modes in a scene, we can combine them into a single program that predicts both modes concurrently. We do this by creating a new tree with an or node as its root. The two programs are its children. This process is repeatable for any number of candidate programs as the newly combined program is now considered a single entity which we can combine with another program. For every possible orientation our language can represent we choose the candidate program with the largest mask. These programs are combined together to produce the final program. Figure 6 shows an example of this process.
Figure 6. Given two programs that produce different placement modes, we combine them into a new program with domain specific operations. We create a new tree with an “or” node as its root. The two programs are set as its children.
@@
Figure 7. Our system maintains performance in both scene synthesis and location distribution modeling with as little as 5% of the original training examples present. Other systems degrade in both final scene quality and the consistency of their per object location distribution.
@@
Table 1. We report FID, Category KL Divergence, and Scene Classifier Accuracy of generated scenes. Methods denoted with “best F1” use the training epoch with the highest F1 score on the location distribution metrics. Methods denoted ”recommended” use the recommended training time. Empty rows mean that the recommended training settings also produce the best F1 scores. Our method generates scenes of comparable quality to previous systems. Although our edge attention and self training algorithm improves our system’s ability to model per object location distributions, they do not significantly hurt or improve its scene synthesis capabilities. ATISS performance degrades when using the epochs with the highest F1 scores because they are often early epochs that have not yet memorized object placements. Fastsynth does not suffer from the same phenomenon.
@@
[…]
TODO
7. Evaluation
We demonstrate how our system produces per object location distributions that are both more complete and more accurate than than previous data-driven autoregressive methods. Our system can model these distributions without sacrificing final scene quality. We also show our system’s superior performance in both modeling this location distribution and scene synthesis when data is sparse.
To demonstrate these claims we evaluate our system against Fastsynth (Ritchie et al., 2018) and ATISS (Paschalidou et al., 2021) – two data-driven autoregressive scene synthesis systems – on the 3DFRONT (Fu et al., 2021) dataset. We train individual models for four scene types: bedrooms, libraries, living rooms, and dining rooms. We evaluate two versions of the baseline methods. One version uses the recommended training time and the other uses the training epoch which maximizes performance on our proposed location distribution metric.
We justify our self training algorithm by evaluating our system without it. We also designed a variant of our method’s generative model which uses an augmented version of the regular self attention mechanism. This was motivated by empirical evidence that the model could not correctly attend to spatial relationships of objects in a scene. See the supplemental material for more details.
7.1. Location Distribution
We evaluate a scene synthesis system’s ability to model per object location distributions by measuring the Precision, Recall, and F1 Score between a location mask of possible centroid locations extracted from the method in question and a ground truth mask that comes from human annotators. We gave human annotators a partial scene and object and asked them to mark all the possible places that object can go. More details on this user study are shown in the supplemental material.
We generated 100 partial scenes and objects for each scene type. These partial scenes and objects follow the placement order used during scene generation. Our system outputs 4 masks that each represent a different orientation of the object, so before comparison it is collapsed into a single mask. For methods that output continuous values such as ATISS and FastSynth we generate masks by binarizing the output values with the threshold value that maximizes the F1 score on the eval set. Masks are dilated before comparison to eliminate sensitivity to small discrepancies.
We show performance of our system over the course of self training as well as performance of the baseline methods in figure 9. Our method adds placements modes increasing recall over the course of self training without sacrificing precision. Baseline methods give the highest F1 scores with low thresholds and early training epochs resulting in fuzzy masks that can cover many modes, but are imprecise. Recall drops when using the recommended settings because longer training time results in overfitting to particular object placements.
Qualitative examples of masks extracted from our methods are shown in Figure 8.
7.2. Scene Synthesis
@@
We evaluate our system’s ability to perform scene generation from a given floor plan. Our programs do not automatically determine the category and size of the next object to place. We use the category prediction module from Ritchie et al. (Ritchie et al., 2018) to predict the object category of the next object to place and randomly sample category dimensions from the dataset of objects.
@@
In accordance with previous work we report the FID scores of top down orthographic renderings of final scenes, the object category KL divergence between a set of generated scenes and an eval set from 3DFRONT, alongside the real fake scene classifier accuracy on the two scenes. Object models closest to the bounding box dimensions of each object are chosen from 3D-FRONT (Fu et al., 2021). Images for FID come from orthographic renderings of generated scenes where each object model receives per-class coloring.
@@
We finetune a pre-trained AlexNet (Krizhevsky et al., 2012) to classify these orthographic renderings and report its classification accuracy on a held out test set. Closer to 50% is better since the classifier struggles to differentiate between real generated scenes. Our quantitative results are shown in Table 1 and qualitative shown in Figure 10.
@@
Our method generates scenes of comparable quality to previous systems. Although our edge attention and self training algorithm improves our system’s ability to model per object location distributions, they do not significantly hurt or improve its scene synthesis capabilities. ATISS performance degrades when using the epochs with the highest F1 scores because they are often early epochs that have not yet memorized object placements. Fastsynth does not suffer from the same phenomenon.
7.3. Data sparsity
One advantage our system has is that its performance does not degrade to the same degree when there are less data samples to train on. Previous systems require many example placements to learn complete distributions. They also rely on overfitting to object arrangements to produce high quality final scenes. Our system’s program bootstrapping algorithm can generalize sparse samples to more general placement rules. These rules allow our system to maintain consistent performance even when training data is sparse. Even if our generative model overfits to particular programs, our program executor guarantees continuity of the location distribution and validity of placement.
We show in Figure 7 the effect that less training examples have on the predicted location distribution and show how our method does not show the same degree of degradation as other methods. We also report the FID score to measure final scene quality with respect to number of samples trained on.
8. Conclusion
In this work, we study how autoregressive indoor scene synthesis systems overfit to object placements seen during training, producing incomplete location distributions. We introduce a new approach to indoor scene synthesis which uses programs as an intermediate representation to address this issue. Our evaluation demonstrates how final scene quality and completeness of the intermediate location distributions used to produce those scenes are often independent. Current scene quality metrics do not accurately reflect this loss in placement diversity. We hope this work brings greater attention to this phenomenon, as well as neurosymbolic methods as a whole.
9. Ethics Statement
Acknowledgements.
Appendix A Further Details on Constraints

### v2

@@ 5.2 Object Encoding
The object encoder encodes object bounding boxes into an embedding vector $\in\mathbb{R}^{d}$. A learned embedding of the object category is concatenated to the raw values of the other attributes and passed through an MLP. Encoding the floor plan, and implicitly the walls, as a single feature vector causes the model to struggle with algebraic quantities such as object to wall distances. This is especially true for floor plans with non-convex geometry. Instead, we encode each wall segment as its own object.
5.3 Program Encoding and Decoding
@@
We represent programs as two separate sequences. The first sequence is the program’s tree structure flattened with prefix notation and embedded with per token learnable embeddings. The second sequence are the constraint attributes concatenated following inorder traversal. Each constraint takes the form (constraint type, query object index, reference object index, direction) or $(c_{j},q_{j},r_{j},d_{j})$. The constraint type and direction receive per-token learnable embeddings. Tokens which represent the query or reference object index use their respective object embeddings generated by the object encoder.
@@
For tokens of fixed vocabulary length such as program structure, constraint type, and direction, an MLP head is enough to decode their tokens. The reference object index however has a variable length vocabulary. The number of objects in a scene is varies. To address this problem, we pass each reference object head through an MLP to form a pointer embedding $v_{j}\in\mathbb{R}^{d}$ [40]. For a matrix of object embeddings $X\in\mathbb{R}^{t\times d}$, where $t$ is the number of objects in the scene, we compute the reference object index $r_{j}$ as
$r_{j}=\text{argmax}(\text{Softmax}(Xv_{j}))$ (1)
@@
The dot product of the pointer embedding with the object embeddings forms a probability distribution over the objects. The reference object is the object with the highest probability mass.
5.4 Alternative Approaches
Using an LLM to generate DSL programs is another possibility, especially given that our work and many existing zero-shot methods both use constraint-based DSLs  [47, 2, 34, 17]. We find however that few-shot prompting GPT-5[24] to generate programs in our DSL results in hallucinated constraints and only similar performance to our data-driven method, as we will show in Section 7. Another option could be fine-tuning a LLM at each iteration of our self-training procedure. However, this approach is computationally infeasible. Fine tuning Llama2-7b [38] with LoRA [15] results in a 515 % increase in training time.
6 Program Self-Training
@@
In this section, we describe our program self-training algorithm which improves next-object distribution locations predicted by our system. An overview of our algorithm is shown in Figure 2.
@@
Our algorithm falls under the PLAD [18] family, a conceptual framework for unsupervised program bootstrapping. These methods iteratively improve a dataset of programs by searching for new and better programs, retraining on them, and then repeating the process. Our algorithm also takes inspiration from Talton et al. [36], a work which uses probabilistic context free grammars (PCFGs) to learn a procedural model from a set of examples. Their optimization begins with the “most specific” grammar and converges to a grammar which is not too specific and not too general.
@@
Our optimization begins with programs that specify a single valid placement. For each object in a scene, we use geometric heuristics to apply every possible constraint to the object so that the extracted program will only place the object where it was originally found. Our iterative self-training pipeline then adds additional placement modes to these “most restrictive” programs through a search and filtering process. Similar to Ganeshan et al. [11] we use domain specific operations to edit a subset of programs per iteration of self training. This is both for computational feasibility and distributional stability between iterations.
6.1 Candidate Program Generation
@@
For a given partial scene and object to add, we search for new programs by sampling the generative model described in Section 5 and then relaxing both the inferred and original program. Program relaxation involves randomly removing constraints from the program tree to produce a new program. This can help generalize the overly restrictive programs produced by our initial naive approach. For example, an object found in the corner of a room might initially be constrained to rest against both walls. Removing one of these constraints would allow for more general placements along one of the walls. Programs which predict no valid placements or a placement anywhere in the scene are removed from consideration.
@@
Each candidate program predicts only one valid orientation. Programs whose predictions cover multiple orientations are split into subtrees that each predict only one. This constraint reduces the number of repeated subtrees present in the final programs and improves quantitative performance of this algorithm. It is worth noting that this constraint does not restrict the possible orientations the final program can represent. Our later program combination step (Section 6.3) produces programs which predict valid placements for multiple different object orientations.
6.2 Object Placement Classifier
@@
Relaxed programs can produce invalid placements. To prevent these programs from entering the training set, we need to filter out bad programs automatically. Hard positives and negatives for programs are hard to generate, so we train a real fake classifier to predict how in or out of distribution an object placement is. Positive examples are generated by randomly subsampling scenes. Negative examples come from randomly perturbing the rotation and location of a single object. This process can generate false negatives (e.g. perturbing a wardrobe such that it rests against the same wall), but in practice it is sufficient for learning a useful decision boundary. False negatives are rare compared to true negatives. We sample a program’s predicted mask multiple times, insert the object in question at those sampled locations, and then compute the object placement classifier’s real probability for each insertion. The program’s final score is the average of these probabilities.
Our classifier architecture uses the object encoder described in Section 5 with an additional learned vector embedding added to the query object vector encoding.
6.3 Combining Programs
@@
Recall that our programs takes on the structure of a CSG tree. If two candidate programs predict two different placement modes in a scene, we can combine them into a single program that predicts both modes concurrently. We do this by creating a new tree with an or node as its root; the two programs are its children. This process is repeatable for any number of candidate programs, as the newly combined program is now considered a single entity which we can combine with another program. For every possible orientation our language can represent, we choose the candidate program with the largest mask. These programs are combined together to produce the final program.
Figure 5: We show the F1 score, precision, and recall of per-object location distributions compared against human annotated masks. The x axis is each iteration of our program bootstrapping algorithm. Baseline methods are flat because they do not use self training. Methods denoted with ”recommended” use the recommended training time and best performing threshold value. Methods denoted with ”best f1” use the settings which maximize F1 score.
Figure 6: Visualization of annotated masks alongside the location distributions predicted by our method, ATISS, Fastsynth, and GPT-5. Green boxes indicate the original placement of the object in the scene. Masks from our method come from the last iteration of self training. The ”recommended” set of masks for Fastsynth and ATISS use the recommended training time and best performing threshold. The ”best f1” set use the settings which maximize F1 scores. Masks for GPT-5 are produced by few-shot prompting the model to generate our DSL programs.
@@
Figure 7: Our system maintains performance with as little as 5% of the original training data. Other systems degrade in the consistency of their predicted per-object location distribution. With less training data recall decreases and precision increases for baseline methods as they overfit to placement locations.
Figure 8: Visualizations of how less training data affects the predicted location distributions of our method, ATISS, and Fastsynth. As data becomes sparse, baseline methods collapse to the placement locations seen during training. Our method predicts a variety of placement locations with as little as 5% of the original training data. It drops only one placement mode in these examples.
7 Evaluation
We demonstrate how our system produces per object location distributions that are both more complete and more accurate than than previous methods. We also show our system’s superior performance in modeling this location distribution when data is sparse.
Our data-driven baselines are Fastsynth [30] and ATISS [26] trained on the 3D-FRONT [8] dataset. An LLM-based method for object location suggestion does not exist, so we few-shot prompt GPT-5-thinking with chain-of-thought  [24]. The closest systems do full scene synthesis  [47, 2, 34, 17], but none have submodules which output location distributions for each object, so we cannot compare against them. We also justify our self training algorithm by evaluating our system without it.
For Fastsynth and ATISS, we train individual models for four scene types: bedrooms, libraries, living rooms, and dining rooms. We evaluate two versions of each model. One version uses the training time recommended by the original authors, and the other uses the training epoch which maximizes performance on our proposed location distribution metric.
Our other baseline is GPT-5-thinking few-shot prompted with chain-of-thought to generate our DSL from a partial scene and object. Scene geometry, furniture in the scene, and the query object are converted to a text representation by listing each object and their attributes. DSL programs are converted to text by writing their tree structure with prefix notation and listing the constraint attributes of each leaf node.
We evaluate two approaches of sourcing which programs to provide as in-context examples in the prompt. Our method receives programs from a heuristically extracted dataset, so the first approach gives GPT-5 the same kind of program data. The example programs in these prompts place objects of the same category as the target object. These programs are flawed however as they are single placement programs that do not benefit from our self-training approach. To measure the upper end of GPT-5 performance, we give the model access to ground truth programs by hand annotating 5 programs for each scene type. GPT-5 performance is expected to improve with access to ground truth.
7.1 Comparing Predicted Location Distribution to Human Annotation
We evaluate a system’s ability to model per object location distributions by measuring the Precision, Recall, and F1 Score between a location mask of possible centroid locations predicted by the method in question and a ground truth-mask that comes from human annotators.
We generated 100 partial scenes and objects for each scene type. Following previous work  [30] these partial scenes and objects follow placement order based an object’s size and frequency within the dataset. Since our programs output four masks (each representing a different orientation of the object), we collapse them into a single mask for comparison to other methods. For methods that output continuous values, such as ATISS and FastSynth, we generate masks by binarizing the output values with the threshold value that maximizes the F1 score on our validation set. Masks are slightly dilated before comparison to eliminate sensitivity to small discrepancies.
We show performance of our system over the course of self training as well as performance of the baseline methods in Figure 5. Our bootstrapping method adds placements modes, increasing recall without sacrificing precision. Baseline methods with high recall have correspondingly low precision because the threshold and training epoch which maximize F1 scores produces distributions that can cover many modes, but are fuzzy and imprecise. Qualitative examples in Figure 6 show how these distributions predict erroneous locations which often cover a majority of the room layout. Recall drops for these methods when they use the recommended settings because longer training times result in memorized object locations. Note how the data-driven baselines in Figure 6 only predict portions of the annotated placement rules, often collapsing to single locations. Our method by contrast predicts a variety of possible placements.
We obtain similar performance to GPT-5 despite having an orders of magnitude smaller model and no ground truth programs to train on. The exact size of GPT-5 is not publicly available, but close open source equivalents  [13] contain hundreds of billions of parameters. Our model has only 2 million. GPT-5 without access to ground truth programs underperforms our method and tends to generate syntactically-incorrect programs or programs with hallucinated constraints (e.g. ”reachable_by_leg”, ”center”, ”front”, ”horizontal”). When syntactically valid programs are executed, they often result in no valid placements, as shown in Figure 6. GPT-5 with access to ground truth programs achieves similar performance to our method, which requires no annotation. Our method can miss placement modes, but our results demonstrate significant improvement in modeling object location distributions when only given access to single location samples.
7.2 Measuring Performance With Less Scene Data
Our system maintains consistent performance when there are fewer data samples to train on. Previous systems degrade in quality because they require many example placements to learn complete distributions. Our system’s program bootstrapping algorithm can generalize sparse samples to more general placement rules. We show with quantitative metrics in Figure 7 and qualitative examples in Figure 8 the effect that fewer training examples have on the predicted location distribution. As data becomes sparse, baseline methods collapse to the placement locations seen during training. Our method predicts a variety of placement location and only ever drops one placement mode in the shown examples.
8 Conclusion
In this work, we study indoor scene object placement. Our approach uses programs and iterative self training to help address the issue of incomplete location distributions. Our evaluation procedure is the first of its kind, and quantifies the performance of object location prediction methods.
9 Ethics Statement
Acknowledgements
Appendix A Scene Synthesis
@@
Table 1: We report FID, Category KL Divergence, and Scene Classifier Accuracy of generated scenes. Methods denoted with “best F1” use the training epoch with the highest F1 score on the location distribution metrics. Methods denoted ”recommended” use the recommended training time. Empty rows mean that the recommended training settings also produce the best F1 scores. Our method generates scenes of comparable quality to previous systems. Although our self training algorithm improves our system’s ability to model per object location distributions, they do not significantly hurt or improve its scene synthesis capabilities. ATISS performance degrades when using the epochs with the highest F1 scores because they are often early epochs that have not yet memorized object placements. Fastsynth does not suffer from the same phenomenon.
@@
Figure 10: Examples of scenes generated by ATISS, Fastsynth, and our method. Our method is capable of generating scenes of comparable quality to previous methods. Note the degradation in final scene quality when ATISS uses the training settings which produce the highest F1 score.
Figure 11: Our system maintains scene synthesis performance even as the number of training examples drops to as little as 5% of the original training set.
@@
We evaluate our system’s ability to perform scene generation from a given floor plan. Our programs do not automatically determine the category and size of the next object to place. We use the category prediction module from Ritchie et al. [30] to predict the object category of the next object to place and randomly sample category dimensions from the dataset of objects.
@@
In accordance with previous work we report the FID scores [14] of top down orthographic renderings of final scenes, the object category KL divergence between a set of generated scenes and an eval set from 3DFRONT, alongside the real fake object placement classifier accuracy on the two scenes. Object models closest to the bounding box dimensions of each object are chosen from 3D-FRONT [8]. Images for FID come from orthographic renderings of generated scenes where each object model receives per-class coloring.
@@
We finetune a pre-trained AlexNet [20] to classify these orthographic renderings and report its classification accuracy on a held out test set. 50% scene classification accuracy indicates that the classifier cannot differentiate between generated and ground truth scenes, so closer to 50% is better. Our quantitative results are shown in Table 1 and qualitative shown in Figure 10.
@@
Our method generates scenes of comparable quality to previous systems. Although our self training algorithm improves our system’s ability to model per object location distributions, they do not significantly hurt or improve its scene synthesis capabilities. ATISS performance degrades when using the epochs with the highest F1 scores because they are often early epochs that have not yet memorized object placements. Fastsynth does not suffer from the same phenomenon.
Section 7.2 demonstrated how our system maintains its ability to model per object location distributions while other methods degrade when there are less data samples to train on. We also show this to be true for scene synthesis and graph FID and SCA scores with respect to the number of data samples trained on. Figure 11 shows these results.
Figures 19 and 20 show more examples of scenes generated by our method and baseline methods
Appendix B Further Details on Constraints